# Theory and Atomistic Simulation of Electrodeposition

### v1

@@ 3.3.1 The Electronic Structure Method
Machine-learned interatomic potentials (MLIPs) are another methodology that can offer high computational efficiency and perform calculations at an ab initio-level of accuracy, if the appropriate training data is supplied. Currently, most MLIPs do not sufficiently capture long-range electrostatics and thus will likely not suffice for electrochemical simulations, which require long-range interactions to also be described, though some studies have successfully captured long-range effects within short-ranged MLIPs153, 154, 155, 156, 157, 158, 159, 160, 161. It should be noted though, that while there exists no universally best MLIP for every problem (as any pair of optimization algorithms should be equivalent when their performances are averaged over all possible problems 220), for practical purposes, care must be taken to ensure the atomic environment is appropriately represented by the machine learning architecture.
Quantum mechanical methods can be used to elucidate many key interactions and processes in electrodeposition. For example, DFT (or an alternative method) can be used to conduct single-point energy calculations or geometry optimizations, the latter resulting in a structure that is an energetic minimum on the potential energy surface. It should be noted that the relaxed structure could be a local minimum rather than a global minimum; various algorithmic processes exist that can be used in conjunction with quantum mechanical methods to identify global and local minima 221, 222, 223. Quantum mechanical methods such as DFT and even semi-empirical methods 224 typically also permit the calculation of the electronic density of states, which can be used to elucidate various interactions. For example, if a copper cation is a few angstroms away from the cathode surface, its projected electronic density of states will be centered at the Fermi level and will contain one electron in the 4 orbital; if it has been deposited, it will have a much broader density of states 89. Such interactions can also be extracted from some machine learning frameworks, which have been developed to predict the electronic properties; such approaches retain full access to the electronic structure at forcefield-level efficiency while capturing QM in an analytically differentiable representation 225.
3.3.2 Structural Model of the Electrode Surface
@@
For homogeneous electrode surfaces which are atomically flat, periodic boundary conditions can be used as an effective method to model extended surfaces and interfaces 197. Here, the surface can be represented as a repeated slab, which is trivially defined using a unit cell that is infinitely repeated in three directions. For a surface, it is important to ensure the unit cell possesses a large enough vacuum in the direction to avoid periodic images from interacting with each other 197. The unit cell should also be large enough to ensure calculations do not suffer from finite size effects 197; this can be assessed by performing convergence tests on different unit cell sizes. The reader is directed to Hofmann et al. 197 for a more detailed review on the repeated slab approach and practical considerations.
@@
However, the surfaces of many electrode materials, particularly semiconducting electrodes, often possess a high degree of heterogeneity and can include structural defects such as point defects, dopants, dislocations, and lower coverages. In fact, defect sites within electrode surfaces have been shown to anchor and stabilize metal nanostructures 108, 226. This stronger interaction between the nanostructure and the defect site can increase the reactivity exhibited by the supported nanostructure 227, 228, 229, 230, 231 in catalytic reactions. Atomistic simulations thus need to be able to account for the heterogeneity of such electrode surfaces. Modeling isolated defects at extended surfaces can be challenging with periodic boundary conditions, as isolated defects within the unit cell acquire a periodicity which can be unphysical. Furthermore, some doped electrode materials typically possess a relatively low dopant concentration e.g. the dopant concentration within boron-doped diamond electrodes is typically around 0.1% 232. To model such an electrode surface using periodic boundary conditions would necessitate a very large unit cell, which can be computationally intractable with higher-level theories.
@@
The challenges associated with the periodic representation of defects can be overcome by creating truncated cluster models. However, this removes the long-range properties of any bulk material and such calculations can be plagued by spurious finite size effects. Embedded cluster calculations are a viable alternative to periodic slab calculations as they acknowledge the intrinsic locality of surface defect chemistry and permit isolated defects to be modeled whilst breaking translational periodicity. Embedded cluster models have been treated using a variety of approaches, such as QM/QM 233, MM/MM and QM/MM, though care must be taken to ensure the embedded region is truncated appropriately, and the interface between the embedded and embedding regions is treated correctly. Such embedded cluster models are also typically computationally cheaper, which makes the application of higher-level theories more straightforward.
4 Simulation of Electrodeposition Processes
@@
4.1 Electrodiffusion of Solvated Ions
@@
An alternative to MD are Monte Carlo simulations, which randomly sample potential configurations of a system rather than solve Newton’s equations of motions. Monte Carlo simulations can therefore be used to gain insights into the probabilistic nature of metal cation motion, as well as other quantities such as the distribution of electrodeposited metal atoms. However, they cannot directly extract dynamical observables connected to time-correlated functions, such as diffusion rates. Kinetic Monte Carlo 235 simulations analyze the time evolution of a system based on the rates of different elementary processes. In fact, kinetic Monte Carlo simulations are particularly useful for studying rare events and processes that occur over large timescales, which are typically not computationally feasible with conventional MD or DFT, making them very applicable for the description of diffusion phenomena during electrodeposition 236, 237, 238, 239, 240.
4.2 Electron Transfer Reactions
@@
The Newns-Anderson model Hamiltonian 90, 91 is a very popular framework that has been used to describe electronic interactions during atomistic simulations of electrodeposition. This framework reduces the complexity of the system while capturing the essential components of electron transfer. However, there are several limitations of the Newns-Anderson model Hamiltonian, such as the absence of electron correlation and the assumption that the electronic coupling between the electrode and the electrolyte is constant. Solvent effects are also not accounted for, rendering the Newns-Anderson Hamiltonian inappropriate for describing electron transfer processes in highly polarized solvents, and the adiabatic approximation is assumed to be valid, where atomic positions are assumed to not instantaneously change with electronic states during electron transfer, which is inappropriate for describing non-adiabatic effects such as coupled electron-proton transfer or dynamically induced electron transfer.
Extensions to the Newns-Anderson model Hamiltonian have been developed to explicitly include key microscopic parameters in a single formulation 97, 241, 242, 243, 244. For example, inspired by Marcus theory 87, the Schmickler-Newns-Anderson Hamiltonian 242, 241 was developed to describe electrochemical electron transfer by incorporating the effect of the solvent. Recently, the Schmickler-Newns-Anderson Hamiltonian has been further modified to account for the electrostatic interaction between the electrode and the redox couple 245, 246, 95 and is expressed as the sum of four terms:
@@
where is the electronic energy of the redox couple, is the distance between the metal cation and the electrode, is the operator for the occupation number of the redox orbital, and is the interaction parameter that characterizes the strength of electronic interactions 245, 95. The last term in Equation (16) accounts for electron transfer between the metal cation and the electrode surface, with and denoting creation and annihilation operators, respectively 245. As the metal cation approaches the electrode surface, will shift towards the Fermi level of the electrode 245, 244, 247. This phenomenon is referred to as Fermi level pinning and occurs due to the stabilization of the cation by the electric field of the cathode and the interaction with the electronic states of the electrode.
@@
where represents a repulsive Coulombic interaction between the redox couple center and the electrode and thus possesses an opposite sign to the electrostatic potential obtained from electronic structure calculations, where the electrostatic potential is calculated from the perspective of electrons 95. The right-hand side of Equation (19) reflects the change in Coulombic interaction due to reduction and as increases, the stabilization due to reduction increases.
@@
The extended Hamiltonian in Equation (15) contains information about both pre-exponential factor and the free energy barrier in Equation (8), and can thus be used to calculate both adiabatic and non-adiabatic OS-ET rates 241. As atomistic simulations can be used to obtain all the quantities that enter Equation (15), it becomes much easier to gain a physical and chemical understanding of OS-ET kinetics.
4.3 Electronucleation
Simulating the amalgamation of surface-adsorbed metal atoms to form metal nanostructures is computationally expensive due to the large number of degrees of freedom and the sheer number of nucleation pathways possible. Some studies investigate the atom-by-atom growth of nanostructures via single-atom absorption, but as mentioned above, this does not hold in reality as entire clusters can merge or fragment. However, analyzing how adsorption energies, cohesive energies, and interatomic distances change as a function of nanostructure size can still provide many valuable insights 228, 248.
@@ 4.3 Electronucleation
However, for a given atom count, metal nanostructures can exhibit numerous metastable geometries. Experimentally, the determination of the geometric ground state of surface-adsorbed nanoclusters is very difficult, and simulations are much more promising for providing insights. However, the reliable identification and optimization of surface-adsorbed metal nanostructures is particularly challenging due to the structural complexity and the large number of degrees of freedom, such as the number of possible metastable geometries, adsorption site, and surface coverage 197. Algorithmic approaches such as basin- 221, 222 and minima-hopping 223 can be used to identify global and local minima. Recently, nested sampling has been extended to calculate coverage-temperature adsorbate phase diagrams by incorporating all relevant configurational contributions to the free energy 249. However, an additional complexity is introduced when attempting to realize the configuration under certain electrochemical conditions.
Isaev et al. 132 sought to study the time dependence of the radius of a single hemispherical nanocluster electrodeposited onto an electrode. In particular, they proposed models of formation and diffusion-controlled growth of a new-phase nanocluster for potentiostatic electrodeposition, cyclic voltammetry, and galvanostatic electrodeposition. The nanocluster growth rate during galvanostatic deposition was found to be much lower than under potentiostatic condition due to the drop in the overpotential that occurs after formation of the nanocluster. However, at larger currents, multiple nucleation will occur and increasing the concentration of the depositing cations would cause the number of clusters to decrease and each cluster size to increase.
5 Capabilities of Software Packages
In order to simulate electrodeposition, users should seek to use established software packages that can account for key electrochemical variables such as solvation effects. The choice of software should be influenced by the task at hand. A good software package should be well-documented and easy to use. Furthermore, softwares should possess efficient computational scalability across multiple cores as atomistic simulations of complex processes such as electrodeposition can be computationally expensive. To this end, several software packages have impressive parallelism of calculations over central processing units (CPUs) and have started to support acceleration of calculations with graphical processing units (GPUs) 250, 251, 252, 253, 254, which can be useful for improving the computational tractability of large systems and complex simulations. Many of the software packages mentioned below are also currently undergoing development to enable integration with machine learning models. This section seeks to provide a non-exhaustive list of software packages for atomistically simulating the various aspects of metal electrodeposition. For a more beginners’ introduction to computational modeling techniques, see Ismail et al. 255.
While there are numerous software packages that contain DFT implementations; David Sherrill et al. 256 is a useful resource that lists the several electronic structure packages that exist. Solid-state codes for condensed-phase systems are generally more applicable than molecular codes. Examples of such codes include (in alphabetical order): the wavelet code BigDFT 257, the plane-wave pseudopotential code CASTEP 258, CP2K 259, the all-electron numeric atomic orbital code FHI-aims 260, the Python-based projector augmented-wave code GPAW 261, 262, 254, NWChem 263, 264, the linear-scaling code ONETEP 265, the Python-based PySCF 266, 267, Quantum ESPRESSO 268, 269, TURBOMOLE 270, and VASP 271. There are many factors to consider when selecting which DFT software package to use. One important choice is the type of basis set, used to represent the Kohn-Sham orbitals. Plane wave approaches rely on delocalized basis functions but require a special treatment for the core electrons, which can be treated using pseudopotentials or projector-augmented waves 197. In contrast, atom-centered basis functions are centered on the nuclei and more natively describe core electrons 197. However, it should be noted that simulations employing plane wave and atom-centered basis functions should give the same results in the converged limit 272. If tight-binding methods are desired in preference to DFT, then software packages such as DFTB+ 273 and xTB 218, 219 can be used to perform atomistic simulations, though care must be taken when selecting parameter sets.
Most DFT software packages include a wide range of density-functional approximations that can be used for calculations, with many of them having interfaces to the LIBXC 199 library of functionals. In addition, most softwares also include dispersion corrections, with many of them interfacing with standalone programs such as DFT-D4 274 or libraries such as libMBD 212 or libvdwxc 275. Many packages now include implementations of implicit solvation models that can be combined alongside QM methods such as DFT, which is key for simulating electrodeposition.
Quantum mechanical software packages that allow for a grand-canonical treatment of the system include CP2K 259, GPAW 262, 261, 254, Quantum ESPRESSO 268, 269, and VASP 271, among others. DFT calculators supported within the popular Python-based ASE 276 can also be turned into GC-DFT calculators by changing only around 20 lines of code, as shown by Melander 143. ASE itself provides interfaces to many software packages that can be used as calculators for many purposes such as energy calculations, geometry optimizations, and other complex algorithmic processes such as nudged elastic band for transition path calculations and global structure searches 276. Recently, the Python-based matscipy package was published, which is interoperable with ASE and includes a matscipy.electrochemistry module to permit electrochemical simulations that describe the motion and spatial distribution of charged species within an external electric field 277. The matscipy.electrochemistry module includes tools that sample discrete coordinate sets from continuum fields and apply steric corrections to avoid any overlap between species 277. It can therefore be used to generate molecular coordinates that resemble ionic distributions within an electrochemical double layer which can be used as the initial configuration for atomistic simulations.
Many software packages also have the ability to interface with others, which can be particularly useful for embedded cluster calculations, for example. Softwares such as the aforementioned Python-based ASE 276 and ChemShell 278 implement a modular approach to embedded cluster models, where a variety of QM and MM codes can be called upon to evaluate the energies and gradients of the various partitioned regions. ChemShell, scriptable using either the Tcl or Python programming languages 278, can call upon numerous QM and MM codes. The recent Atomic Simulation Interface (ASI), which has a native C-style API, also defines a set of functions that support both classical and ab initio MD, and hybrid QM/MM calculations 279. ASI provides an efficient way to import and export large arrays from electronic structure codes, such as the Hamiltonian, overlap, and density matrices, that are typically monolithic and can be used to develop universal and interoperable codes without sacrificing efficiency for portability 279.
Large-scale molecular mechanical modeling of materials are most efficiently achieved with established softwares such as DL_POLY 280, the Python-based HOOMD-blue 281, and LAMMPS 282. LAMMPS in particular is well-equipped to conduct electrochemical simulations, comprising implementations of packages such as ELECTRODE 283, which enable computationally efficient simulations of electrochemical reactions, electrode potentials, and ion transport within an electrochemical cell. MD simulations have become increasingly popular as they are computationally well-suited to capitalize upon the hardware speedups bestowed by Moore’s law and the resulting large-scale parallelism over both CPUs and GPUs in recent years 282. Efficient MD codes can therefore be used to simulate systems from atomic to mesoscale length scales at picosecond to microsecond timescales 282. Within MD, the computational cost for models with short-range interactions has formal linear scaling i.e. , and long-range Coulombic interactions have formal scaling 282.
6 Applications and Challenges
New experimental techniques are emerging that can provide valuable data to support and validate atomistic simulations. For example, electron microscopy methods have been a popular choice to study metal electrodeposition due to the high resolution they offer 284, 285, 286, 287, 26. Scanning transmission electron microscopy has been shown to be capable of dynamically visualizing the early stages of electronucleation for metals such as gold, with structural resolution on the atomic scale and time resolution defined by the sequential analysis of short electrodeposition runs (several milliseconds of deposition per run) 26. Transmission electron microscopy has also been used to study the electrodeposition of metals such as lithium, silver, nickel and platinum at submicroscopic resolution 284, 285, 288, 286, 287, 289. Several studies also report the use of scanning electron microscopy to investigate metal electrodeposition 290, 286, 291. While liquid-cell transmission electron microscopy has made lots of progress in monitoring dynamic electrochemical systems 292, 293, it has limited resolution due to factors such as electron-beam-induced gas bubble formation and electron scattering in the liquid 26; in contrast, ex situ aberration-corrected scanning transmission electron microscopy is not only capable of resolving single atoms but can be used to quantify the number of atoms within a particle 294, 26.
In contrast, scanning probe methods such as scanning tunnelling microscopy and atomic force microscopy generate images of surfaces using a physical probe that scans the sample 295. Both atomic force microscopy 296 and scanning tunnelling microscopy 297, 298, 299, 300, 301, 302, 303 are capable of analyzing the influence of current on the electrodeposited structure at submicroscopic resolutions. Scanning electrochemical cell microscopy has also been used to study the initial electronucleation stages and mobility of metals such as platinum 304, copper 305, iron 306 and silver 25, and has gained much attention 307, 308, 309 due to its ability to routinely operate at submicroscopic scales 305, 310, 311. Other microscopy techniques such as surface plasmon resonance microscopy 312 and dark-field scattering microscopy 313, 314 also exist and have been used to investigate metal electrodeposition. Surface plasmon resonance microscopy is a non-intrusive optical technique that is highly sensitive to nanoscopic objects, while dark-field scattering microscopy has a relatively simple experimental setup that can directly probe the plasmonic properties of individual structures and allow for correlation with electron microscopy 313, 314. Their performance, however, is restricted by the very small field of view which increases the difficulty in acquiring quantitative data on electronucleation 312, 313. Wide-field surface plasmon resonance microscopy, however, removes this constraint and allows for the growth of hundreds of nuclei to be tracked simultaneously at a reasonable time resolution () 312.
Advancements in the theory and atomistic simulation of metal electrodeposition have helped to elucidate many phenomena. However, several open challenges remain that hinder the full realization of accurate models. One primary challenge is the precise modeling of the electrochemical interface. The dynamic nature of this interface requires sophisticated simulation methods capable of capturing these nuances. Additionally, incorporating quantum mechanical effects into large-scale simulations without incurring prohibitive computational costs is an ongoing challenge.
To model electrodeposition, the electrosorption, electron transfer, on-surface migration and nucleation processes need to be captured. So far, no simulation approach exists that captures all these aspects atomistically. Bridging the gap between the femtosecond timescales of electron transfer and atomic motions, and the multisecond timescales of electrodeposition processes demands the development of advanced algorithmic approaches. Understanding electronucleation and growth mechanisms at the atomic level, particularly the role of surface defects and heterogeneities, pose additional challenges. The effects of solvent molecules and ions, including solvation dynamics and ion pairing, further complicate these simulations. Continuum and atomistic approaches need to be combined to create multiscale approaches that can be to used to tackle such challenges. It is evident that MLIPs and other ML surrogate models coupled with electronic structure theory 315 will have to play a crucial role in developing a unified multi-scale modelling description of electrodeposition.
@@ 6 Applications and Challenges
Both theory and experiment will be needed to answer the open questions in the field. In terms of electronucleation, this includes consideration of non-classical growth pathways such as surface migration, aggregation and coalescence of small nanoclusters 25, 118; formation of metastable clusters into crystalline nanoparticles 26; and nucleation and dissolution events that occur before stable nuclei form 316. Other open questions 317, 111 include why the measured number of nuclei is higher than the calculated number of active sites and why single atoms are so stable. With the emergence of improved simulation methods and experiments with high spatial and temporal resolution, there is promise for these open questions to be answered in the near future.

### v2

@@ Introduction
MLIPs are another methodology that can offer high computational efficiency and perform calculations at an ab initio-level of accuracy, if the appropriate training data is supplied. Currently, most MLIPs do not sufficiently capture long-range electrostatics and thus will likely not suffice for electrochemical simulations, which require long-range interactions to also be described, though some studies have successfully captured long-range effects within short-ranged MLIPs239, 240, 241, 242, 243, 244, 245, 246 and further developments on explicit long-range MLIPs are underway. 247, 248, 249
Structural Model of the Electrode Surface
@@
For homogeneous electrode surfaces which are atomically flat, periodic boundary conditions can be used as an effective method to model extended surfaces and interfaces 216. Here, the surface can be represented as a repeated slab, which is trivially defined using a unit cell that is infinitely repeated in three directions. For a surface, it is important to ensure the unit cell possesses a large enough spacing in the -direction to avoid periodic images from interacting with each other 216. The unit cell should also be large enough to ensure calculations do not suffer from finite size effects 216; this can be assessed by performing convergence tests on different unit cell sizes. The reader is directed to Hofmann et al. 216 for a more detailed review on the repeated slab approach and practical considerations.
@@
However, the surfaces of many electrode materials, particularly semiconducting electrodes, often possess a high degree of heterogeneity and can include structural defects such as point defects, dopants, dislocations, and lower coverages. In fact, defect sites within electrode surfaces have been shown to anchor and stabilize metal nanostructures 125, 250. This stronger interaction between the nanostructure and the defect site can increase the reactivity exhibited by the supported nanostructure 251, 252, 253, 254, 255 in catalytic reactions. Atomistic simulations thus need to be able to account for the heterogeneity of such electrode surfaces. Modeling isolated defects at extended surfaces can be challenging with periodic boundary conditions, as isolated defects within the unit cell acquire a periodicity which can be unphysical. Furthermore, some doped electrode materials typically possess a relatively low dopant concentration e.g. the dopant concentration within boron-doped diamond electrodes is typically around 0.1% 256. To model such an electrode surface using periodic boundary conditions would necessitate a very large unit cell, which can be computationally intractable with higher-level theories.
@@
The challenges associated with the periodic representation of defects can be overcome by creating truncated cluster models. However, this removes the long-range properties of any bulk material and such calculations can be plagued by spurious finite size effects. Embedded cluster calculations are a viable alternative to periodic slab calculations as they acknowledge the intrinsic locality of surface defect chemistry and permit isolated defects to be modeled whilst breaking translational periodicity. Figure 5 shows how electrode surfaces can be represented within atomistic simulations depending on which region of the surface is of interest. Embedded cluster models have been treated using a variety of approaches, such as QM/QM 257, MM/MM and QM/MM, though care must be taken to ensure the embedded region is truncated appropriately, and the interface between the embedded and embedding regions is treated correctly. Such embedded cluster models are also typically computationally cheaper, which makes the application of higher-level theories more straightforward.
Simulation of Electrodeposition Processes
@@
Electrodiffusion in the Electrolyte and Mass Transport
@@
To address the limitations of atomistic methods in modeling longer timescales and larger systems, mesoscale simulation techniques such as continuum models 154, 155, 156, 157, 158 and kinetic Monte Carlo 259, 260, 261, 156 play a crucial role in modeling electrochemical reaction conditions and mass transport over larger length and time scales. Continuum models, which typically solve the Poisson-Nernst-Planck equations, provide a macroscopic description of ion transport, charge redistribution, and electrostatic potential gradients within the electrolyte. These models are particularly useful in capturing mass transport effects, including diffusion, migration, and convection, which strongly influence electrodeposition kinetics and morphological evolution. Phase-field methods 262, 263, 264, 265, 266, 267, 268 introduce order parameters to describe interfacial morphology and dynamic structural changes. Unlike continuum models, phase-field approaches do not track individual atoms but rather model the smooth transition between phases. By incorporating thermodynamic and kinetic driving forces, phase-field approaches enable the study of morphological evolution under different electrochemical conditions, complementing continuum models that focus on ion transport and charge redistribution.
Mass transport plays a critical role in electrodeposition, particularly in systems where diffusion limitations lead to concentration gradients that influence dendrite formation. For example, the phase-field model developed by Cogswell 265 incorporates Marcus kinetics to describe charge transfer in concentrated solutions and allows for simulations of dendritic growth under mass transport constraints. This model, formulated in terms of the grand electrochemical potential, enables the widening of interface regions to reach experimentally relevant length- and timescales whilst enforcing electroneutrality. Notably, the model achieves quantitative agreement with experimental data and predicts that reducing the exchange current, which reflects how readily electrons are transferred between the electrode and ions in solution, can suppress dendrite growth. Screening electrolytes based on their exchange currents could thus serve as a strategy for controlling dendritic instability.
Electron Transfer Reactions
@@
The Newns-Anderson model Hamiltonian 269, 270 is a popular framework that has been used to describe electronic interactions during atomistic simulations of electrodeposition. This framework reduces the complexity of the system while capturing the essential components of electron transfer. However, there are several limitations of the Newns-Anderson model Hamiltonian, such as the absence of electron correlation and the assumption that the electronic coupling between the electrode and the electrolyte is constant. Solvent effects are also not accounted for, rendering the Newns-Anderson Hamiltonian inappropriate for describing electron transfer processes in highly polarized solvents. Furthermore, the adiabatic approximation is assumed to be valid, where atomic positions are assumed not to instantaneously change with electronic states during electron transfer, which is inappropriate for describing non-adiabatic effects such as coupled electron-proton transfer or dynamically induced electron transfer.
Extensions to the Newns-Anderson model Hamiltonian have been developed to explicitly include key microscopic parameters in a single formulation 121, 271, 272, 273, 274. For example, inspired by Marcus theory 105, the Schmickler-Newns-Anderson Hamiltonian 272, 271 was developed to describe electrochemical electron transfer by incorporating the effect of the solvent. Recently, the Schmickler-Newns-Anderson Hamiltonian has been modified to account for the electrostatic interaction between the electrode and the redox couple 275, 276, 118 and is expressed as the sum of four terms:
@@
where is the electronic energy of the redox couple, is the distance between the metal cation and the electrode, is the operator for the occupation number of the redox orbital, and is the interaction parameter that characterizes the strength of electronic interactions 275, 118. The last term in Equation (7) accounts for electron transfer between the metal cation and the electrode surface, with and denoting creation and annihilation operators, respectively 275. As the metal cation approaches the electrode surface, will shift towards the Fermi level of the electrode 275, 274, 277. This phenomenon is referred to as Fermi level pinning and occurs due to the stabilization of the cation by the electric field of the cathode and the interaction with the electronic states of the electrode.
@@
where represents a repulsive Coulombic interaction between the redox couple center and the electrode and thus possesses an opposite sign to the electrostatic potential obtained from electronic structure calculations, where the electrostatic potential is calculated from the perspective of electrons 118. The right-hand side of Equation (10) reflects the change in Coulombic interaction due to reduction and as increases, the stabilization due to reduction increases.
@@
The extended Hamiltonian in Equation (6) can be parametrized from first-principles using a selection of the techniques discussed above and used to predict both the pre-exponential factor and the free energy barrier in rate equations, and can thus be used to calculate both adiabatic and non-adiabatic ET rates 271. As atomistic simulations can be used to obtain all the quantities that enter Equation (6), this approach provides physical and chemical understanding of ET kinetics as a function of experimentally accessible parameters.
Schmickler-Newns-Anderson-based models have been a powerful theoretical framework for describing OS-ET at interfaces by successfully capturing key aspects of electronic coupling and reorganization energy 278, 118, 279. Several studies have combined the model with MD to study electrochemical proton-coupled electron transfer at electrode-solution interfaces 280, 281, 278. While this approach provides a rigorous way to connect electronic structure theory with electrochemical kinetics, its direct application to electrodeposition remains largely unexplored. The fundamental principles of the model could, in principle, be extended to describe inner-sphere charge transfer and metal adatom stabilization, few studies have pursued this direction. Given the complexity of electrodeposition, where electron transfer is coupled to electrosorption, surface diffusion, and electronucleation processes, extensions of the Schmickler-Newns-Anderson model to explicitly account for these effects would be highly desirable. Future work should explore how modifications of this framework could provide new insights into the electronic structure of intermediates during the electrodeposition process, and how the electrode potential governs electronucleation rates. Developing such extensions could bridge the gap between existing electronic structure models and experimentally observed electrodeposition kinetics.
Electronucleation and Growth
Simulating the amalgamation of surface-adsorbed metal atoms to form metal nanostructures is computationally expensive due to the large number of degrees of freedom and the sheer number of nucleation pathways possible. Some studies investigate the atom-by-atom growth of nanostructures via single-atom absorption, but as mentioned above, this does not hold in reality as entire clusters can merge or fragment. In this context, it is important to consider the size-dependent surface mobility of nanoclusters, which plays a critical role in determining their likelihood of coalescence or fragmentation. Smaller clusters typically exhibit higher mobility, allowing them to migrate across the electrode surface and contribute to more dynamic growth pathways, whereas larger clusters may become kinetically trapped. These effects can significantly influence both the spatial distribution and the growth kinetics of electrochemical nanostructures. Nevertheless, analyzing how adsorption energies, cohesive energies, and interatomic distances change as a function of nanostructure size can still provide many valuable insights 252, 282. An additional complexity is introduced when attempting to simulate structures and atomic-scale growth under certain electrochemical conditions. Isaev et al. 283 sought to study the growth kinetics of a single hemispherical nanocluster electrodeposited onto an electrode. In particular, they proposed models of formation and diffusion-controlled growth of a new-phase nanocluster for potentiostatic electrodeposition, cyclic voltammetry, and galvanostatic electrodeposition. The nanocluster growth rate during galvanostatic deposition was found to be much lower than under potentiostatic conditions due to the drop in the overpotential that occurs after formation of the nanocluster. However, at larger currents, multiple electronucleation events will occur and increasing the concentration of the depositing cations would cause the number of clusters to decrease and each cluster size to increase.
@@
For a given cluster size, metal nanostructures can exhibit numerous metastable geometries. Experimentally, the determination of the geometric ground state of surface-adsorbed nanoclusters is challenging, and simulations are much more promising for providing insights. However, the reliable identification and optimization of surface-adsorbed metal nanostructures is particularly challenging due to the structural complexity and the large number of degrees of freedom, such as the number of possible metastable geometries, adsorption site, and surface coverage 216. Global structure search approaches such as basin- 284, 285 and minima-hopping 286 can be used to identify stable structures. Recently, nested sampling has been extended to calculate coverage-temperature adsorbate phase diagrams by incorporating all relevant configurational contributions to the free energy 287. Such stable structure models can be directly compared to ex-situ imaging techniques of electrodeposited nanostructures. Combined with sequences of short electrodeposition and ex situ imaging, this can provide insights into the structural evolution as a function of macroscopic time. 26.
To bridge the gap between atomistic resolution and experimentally relevant timescales, mesoscale methods such as phase-field methods and kinetic Monte Carlo can be applied. Phase-field methods 262, 263, 264, 265, 266, 267, 268 model the smooth transition between metal and electrolyte phases using order parameters and incorporate both thermodynamic and kinetic driving forces to simulate morphological evolution. For example, they enable simulations of growth, coarsening, and competitive nucleation under varying electrochemical conditions. Kinetic Monte Carlo methods 259, 260, 261, 156, on the other hand, provide a statistical framework for modeling the stochastic nature of electronucleation and the impact of thermal fluctuations on electrodeposition. They also enable the time evolution of nucleation and growth, based on transition rates between configurations, to be simulated. This is especially valuable for capturing rare events and long-time processes that are inaccessible to standard MD. The combination of atomistic and mesoscale techniques therefore provides a powerful toolkit for understanding electronucleation and growth mechanisms across length and time scales.
Looking forward, mesoscale simulation techniques could complement atomistic simulations by bridging the gap between detailed quantum mechanical descriptions and experimentally observable macroscopic behavior. While mesoscale models sacrifice some atomic detail, their computational efficiency and ability to capture large-scale mass transport effects make them indispensable for a comprehensive understanding of electrodeposition. By integrating continuum, phase-field, and kinetic Monte Carlo approaches with atomistic insights, robust multiscale frameworks could be developed to predict the interplay of ion transport, surface adsorption, and growth. Such a holistic approach would not only enhance theoretical understanding but also facilitate the design of more stable, controlled, and efficient electrodeposition processes by optimizing electrolyte composition and deposition parameters.
Conclusions and Outlook
Advancements in the theory and atomistic simulation of metal electrodeposition have helped to elucidate many phenomena. Current successes include: accurate modeling of OS-ET reaction kinetics using Marcus-type theories 101; continuum simulations that capture mass transport, double-layer formation, and electrostatic potential gradients 157; and identification of the role of surface defects and heterogeneities in nucleation through atomistic simulations 282. However, several open challenges remain that hinder the full realization of holistic models of electrodeposition. For instance, while simulation methods can determine the required overpotential, the exchange current density, and reaction free energies, linking these parameters directly to experimentally measured Tafel slopes, experimentally employed overpotentials and the realistic electrode morphology is still problematic.
There are notable areas that are still work-in-progress, namely: a unified atomistic simulation framework that captures all aspects of the deposition process, from femtosecond electron transfer events to macroscopic growth kinetics, remains elusive. Such a framework would enable the direct prediction of electronucleation rates, surface morphologies, and growth modes under realistic electrochemical conditions, which currently require empirical input or simplified models. The transition from fundamental charge transfer events to macroscopic deposition patterns remains poorly understood, and a lack of a unified framework would significantly hinder the ability to design materials and processes with tailored electrodeposition characteristics. Secondly, the dynamic nature of the electrochemical interface, including the complex interplay between solvent reorganization, ion pairing, and surface diffusion, is still only partially understood. For example, experimentally observed deviations from classical Tafel behavior in metal electrodeposition, such as unexpectedly high or potential-dependent Tafel slopes, suggest that solvent dynamics and ion rearrangements at the interface play a crucial role in charge transfer kinetics. However, these effects are difficult to disentangle experimentally and remain challenging to model accurately at the atomic scale. Finally, the coupling of interfacial ion and electron transfer to mass transport in the electrolyte in a predictive, ab initio manner is still not fully developed.
Long-term challenges include: developing multiscale modeling frameworks that efficiently integrate ab initio simulations with mesoscale simulation methods; achieving quantitative agreement between measured Tafel slopes and transfer coefficients from operando experiments and computational predictions; and designing robust MLIPs that can accurately capture long-range interactions and electrostatic effects in realistic cathode/electrolyte interfaces. It is evident that MLIPs and other ML surrogate models coupled with electronic structure theory 288 will play a crucial role in developing a unified multiscale modeling description of growth and nucleation at electrified interfaces 157.
Going forward, specific actions for the computational electrochemistry community should include: strengthening collaborations with experimental groups to validate and refine simulation models through direct comparison with high-resolution operando data; developing standardized benchmarks for electrodeposition simulations, focusing on key observables such as overpotential, exchange current densities, electronucleation rates, and morphology evolution; advancing algorithms to bridge the time and length-scale gap, particularly through hybrid multiscale methods that integrate ab initio insights with continuum modeling; and focusing on the parameterization of coupled ion-electron transfer theories using ab initio data to better reproduce experimental voltammetry and Tafel slopes.
In summary, while significant progress has been made, a disconnect remains between detailed atomistic simulations and macroscopic experimental observables. Bridging this gap requires a concerted effort to integrate quantum mechanical insights with multiscale modeling approaches, thereby enabling more accurate and predictive simulations of electrodeposition processes. Crucially, to reduce the complexity gap and facilitate experiment-theory synergy, this will, at least initially, require experimentation on heavily simplified systems and idealized model systems. A necessary detour that will provide significant rewards in terms of improved models and understanding. Enhanced collaboration between theorists and experimentalists is essential to refine models, validate simulation outputs, improve the interpretation of experiments, and ultimately achieve a unified description of metal electrodeposition.
@@
Both theory and experiment will be needed to answer the open questions in the field. In terms of electronucleation, this includes consideration of non-classical growth pathways such as surface migration, aggregation and coalescence of small nanoclusters 25, 135, 37; formation of metastable clusters into crystalline nanoparticles 26; and nucleation and dissolution events that occur before stable nuclei form 289, 171. Other open questions 152, 128 include why the measured number of nuclei is higher than the calculated number of active sites and why single atoms are so stable. With the emergence of improved simulation methods and experiments with high spatial and temporal resolution, there is promise for these open questions to be answered in the near future.